# Exploring dynamical quantum phase transition from pure states to mixed states through generalized Su-Schrieffer-Heeger models

### v1

@@ Abstract
We investigate dynamic quantum phase transitions (DQPTs) in both pure and mixed states within the framework of the generalized SSH model, specifically analyzing the SSH-3 and SSH-4 models, which exhibit different symmetries. We find that the SSH-3 model, characterized by a chiral-like point symmetry rather than true chiral symmetry, supports robust localized edge states associated with its topological properties. Our results show that DQPTs for pure states occur following a quench that crosses the topological transition, even with an open energy band gap. For mixed states, DQPT behavior is consistent at low temperatures, but significant changes are observed at high temperatures, resulting in the emergence of multiple critical times. In contrast, the SSH-4 model, which possesses chiral symmetry, allows for the analysis of two distinct energy spectrum configurations. We conclude that the occurrence of DQPTs for pure states in the SSH-4 model necessitates a quench from an initial state without a band gap while crossing the critical point of the topological transition, whereas DQPTs are absent for mixed states at elevated temperatures.
@@ I Introduction
Recent experimental investigations on ultra-cold atoms confined in optical lattices [1, 2] have significantly advanced the exploration of non-equilibrium dynamics in isolated quantum systems [3]. A primary focus within this field is the temporal evolution of a quantum system following a sudden global quench, which can be readily performed in experiments and studied theoretically [4]. In these scenarios, the Loschmidt echo, indicating the overlap between the pre- and post-quench Hamiltonian eigenstates, plays a pivotal role [5]. The formal parallelism between the Loschmidt amplitude and the canonical partition function of an equilibrium system has led to the concept of dynamical quantum phase transitions (DQPTs), aiding in the comprehension of phase and phase transition concepts in non-equilibrium settings [6, 7, 8].
@@
The concept of dynamical quantum phase transitions (DQPTs) elucidates the critical behaviors of the Loschmidt echo, $\mathcal{L}(t)=|\mathcal{G}(t)|^{2}$, in the non-equilibrium evolution of quantum systems. The Loschmidt amplitude, $\mathcal{G}(t)$, quantifying the overlap between evolving and initial states, is expressed as $\mathcal{G}(t)=\langle\psi_{0}|e^{-iHt}|\psi_{0}\rangle$. To reveal DQPTs, the dynamical free energy density as the rate function of the Loschmidt echo in the thermodynamic limit, $\lambda(t)=-\lim_{N\rightarrow+\infty}\frac{1}{N}\ln{[\mathcal{L}(t)]}$, or its time derivative, exhibits cusp-like singularities at critical times. DQPTs have been extensively analyzed in various quantum systems, encompassing XY chains [9, 10, 11, 12], Kitaev honeycomb models [13], non-integrable models [14, 15, 16, 17, 18], systems with long-range interactions [19, 20, 21, 22, 23, 24, 25], quantum Potts models [26], non-Hermitian systems [27, 28, 29, 30], Bose-Einstein condensates [31], periodically driven systems [32, 33, 34, 35, 36, 37, 38, 39], mixed states [40, 41, 42, 43, 44, 45, 46, 47, 48, 49], among others. Experimental validations of DQPTs have been achieved through trapped ions simulations [50, 51, 52, 53], nuclear magnetic resonance quantum simulators [54], quantum walks of photons [55, 56], and spinor condensate simulations [57]. Moreover, an alternative definition of DQPTs examines the late-time behavior of order parameters [58, 59, 60, 61, 62], revealing two types of DQPTs in the long-range quantum Ising chain [24].
Recent advancements have unveiled that DQPTs can shed light on the topological aspects of real-time quench dynamics [63]. In this regard, a bulk topological quantum number, derived from the momentum space winding number of the Pancharatnam geometric phase, has been introduced to encapsulate these topological characteristics, known as the dynamic topological order parameter (DTOP). The DQPTs involving topological properties have received some attention, especially in band models [64, 65, 66]. However, whether existing conclusions can be generalized to quantum band systems with certain special symmetries remains an open question. For example, in the case of SSH-3, despite having robust localized edge states, it lacks the chiral symmetry and topological invariants characteristic of the SSH model [67]. Therefore, in this paper, we focus on the influence of the symmetry of band models on the properties of dynamical quantum phase transitions. To this end, we study the extended SSH model, primarily analyzing the dynamical quantum phase transition behavior of two typical models after quenching. We also discuss the impact of finite temperatures on the dynamical behavior.
@@
where $t_{1},\cdots,t_{m-1}$ represent intracell couplings, and $t_{m}$ denotes the intercell coupling. In contrast to the SSH model, the generalized SSH-m model reveals richer variety of topological phase transitions [67, 68]. In the Hamiltonian (1), we have imposed the periodic boundary condition. Consequently, the Bloch Hamiltonian is obtained by
@@
In contrast, when considering the dynamics at finite temperature, the initial state is prepared as the thermal state described by a full-ranked density matrix $\rho_{0}$. The density matrix at time $t$ after the quench is given by $\rho(t)=e^{-iH(t)}\rho_{0}e^{iHt}$. Hence, a generalization of the Loschmidt amplitude to general density matrices we consider is defined as [41, 40]
@@
It is important to emphasize that the generalized Loschmidt amplitude proposed here is not limited by the parallel transport condition [69, 70, 42, 71]. In our model, the initial density matrices for every $k$ take the form
@@ III DQPTs for pure states in SSH-3 model
In this section, we consider a generalization of the SSH model hosting three lattice sites in a unit cell. Notably, for the 1D SSH-3 model, the spectrum of the bulk Hamiltonian is not symmetric concerning zero [67], implying that the system does not possess chiral symmetry. The winding number is also not well-defined for the SSH-3 model [72]. However, the bulk Hamiltonian exhibits an additional interesting symmetry called point chirality, i.e. $\Gamma_{k}H_{k}\Gamma_{k}^{{\dagger}}=-H_{\pi-k}$ with $\Gamma_{k}=\mathrm{diag}(1,-1,1)$. In Fig. 1, we show the energy spectra of the bulk Hamiltonians for varying $t_{3}$ with fixed inner-cell hopping parameters $t_{1}$ and $t_{2}$. The Hamiltonian exhibits the gap closing only when $t_{1}=t_{2}=t_{3}$, which exactly corresponds to the case of the SSH model.
@@ V DQPTs in SSH-4 model
By calculating the energy spectra under different parameters, we have identified three typical energy spectrum configurations in the SSH-4 model, one of which resembles the energy spectra of the period-two Kitaev chain, as discussed in the Ref. [66]. In this article, we will discuss two other cases different from the period-two Kitaev chains, with the shapes of their energy spectra shown in Fig. 8.
@@
From Fig. 8 (a), it is can be seen the system with parameters $t_{1}=t_{2}=t_{3}$ has a gap-closing point at $t_{4c}=1$. In contrast, as shown in Fig. 8 (b), the system with parameters $t_{1}=t_{3}=1$ and $t_{2}=2$ has two gap-closing point at $t_{4c}=0.5$ and $t_{4c}=2$. These critical points that collapse the energy bands will result in the topological transitions in the system, which are characterized by the change in the number of edge modes [68]. It is evident that based on previous experience, we can speculate that in the SSH-4 model, the behavior of the DQPT after a quench will depend on whether the quench passes through these critical points.

### v2

# Exploring dynamical quantum phase transition from pure states to mixed states through extended Su-Schrieffer-Heeger models

@@ Abstract
We investigate dynamic quantum phase transitions (DQPTs) in both pure and mixed states within the framework of the extended SSH model, specifically analyzing the SSH-3 and SSH-4 models, which exhibit different symmetries. We find that the SSH-3 model, characterized by a chiral-like point symmetry rather than true chiral symmetry, supports robust localized edge states associated with its topological properties. Our results show that DQPTs for pure states occur following a quench that crosses the topological transition, even with an open energy band gap. For mixed states, DQPT behavior is consistent at low temperatures, but significant changes are observed at high temperatures, resulting in the emergence of multiple critical times. In contrast, the SSH-4 model, which possesses chiral symmetry, allows for the analysis of two distinct energy spectrum configurations. We conclude that the occurrence of DQPTs for pure states in the SSH-4 model necessitates a quench from an initial state without a band gap while crossing the critical point of the topological transition, whereas DQPTs are absent for mixed states at elevated temperatures.
@@ I Introduction
Recent experimental investigations on ultra-cold atoms confined in optical lattices Bloch et al. (2008); Lewenstein et al. (2012) have significantly advanced the exploration of non-equilibrium dynamics in isolated quantum systems Polkovnikov et al. (2011). A primary focus within this field is the temporal evolution of a quantum system following a sudden global quench, which can be readily performed in experiments and studied theoretically Mitra (2018). In these scenarios, the Loschmidt echo, indicating the overlap between the pre- and post-quench Hamiltonian eigenstates, plays a pivotal role Zurek et al. (2005). The formal parallelism between the Loschmidt amplitude and the canonical partition function of an equilibrium system has led to the concept of dynamical quantum phase transitions (DQPTs), aiding in the comprehension of phase and phase transition concepts in non-equilibrium settings Heyl et al. (2013); Zvyagin (2016); Heyl (2018).
@@
The concept of dynamical quantum phase transitions (DQPTs) specifically elucidates the critical behaviors of the Loschmidt echo, $\mathcal{L}(t)=|\mathcal{G}(t)|^{2}$, in the non-equilibrium evolution of quantum systems. The Loschmidt amplitude $\mathcal{G}(t)$, quantifying the overlap between evolving and initial states, is expressed as $\mathcal{G}(t)=\langle\psi_{0}|e^{-iHt}|\psi_{0}\rangle$. To manifest DQPTs, the dynamical free energy density is defined as the rate function of the Loschmidt echo in the thermodynamic limit, $\lambda(t)=-\lim_{N\rightarrow+\infty}\frac{1}{N}\ln{[\mathcal{L}(t)]}$, which exhibits cusp-like singularities at critical times. Recent advancements have unveiled that DQPTs can shed light on the topological aspects of real-time quench dynamics Budich and Heyl (2016). The DQPT has been extensively studied in various quantum systems, including integrable models (e.g. XY chains Vajna and Dóra (2014); Cao et al. (2022); Porta et al. (2020), Kitaev honeycomb model Schmitt and Kehrein (2015)), non-integrable models Karrasch and Schuricht (2013); Andraschko and Sirker (2014); Heyl (2014); Kriel et al. (2014); Sharma et al. (2015), systems with long-range interactions Halimeh and Zauner-Stauber (2017); Homrighausen et al. (2017); Obuchi et al. (2017); Zauner-Stauber and Halimeh (2017); Dutta and Dutta (2017); Žunkovič et al. (2018); Halimeh et al. (2020), non-Hermitian systems Zhou et al. (2018); Mondal and Nag (2022); Zhai et al. (2022); Mondal and Nag (2023), inhomogeneous systems Yang et al. (2017); Yin et al. (2018); Srivastav et al. (2019); Cao et al. (2020); Mishra et al. (2020); Modak and Rakshit (2021); Hoyos et al. (2022); Cao et al. (2023); Kuliashov et al. (2023), and among others Karrasch and Schuricht (2017); Lahiri and Bera (2019); Abdi (2019); Huang et al. (2019); Zache et al. (2019); Jafari et al. (2019). Besides the quantum quench, other nonequilibrium protocols, for instance, the linear ramp Divakaran et al. (2016); Sharma et al. (2016); Puskarov and Schuricht (2016); Zamani et al. (2024); Cao et al. (2024a), periodic Floquet systems Yang et al. (2019); Zamani et al. (2020); Shirai et al. (2020); Zhou and Du (2021); Jafari and Akbari (2021); Hamazaki (2021); Zamani et al. (2022); Jafari et al. (2022), and systems subject to noise Jafari et al. (2024), have also attracted wide attention. Moreover, experimental validations of DQPTs have been achieved through trapped ions simulations Fläschner et al. (2017); Jurcevic et al. (2017); Chen et al. (2020); Muniz et al. (2020), nuclear magnetic resonance quantum simulators Nie et al. (2020), quantum walks of photons Wang et al. (2019); Xu et al. (2020), and spinor condensate simulations Tian et al. (2020). Additionally, a similar definition of the dynamical phase transition, studying the asymptotic late-time steady state of the order parameters Barmettler et al. (2009); Eckstein et al. (2009); Sen et al. (2016), also arouses a lot of interest. Two types of DQPTs have been found related to the long-range quantum Ising chain Žunkovič et al. (2018).
Recent studies have intriguingly extended the theory of DQPTs to nonzero temperatures Abeling and Kehrein (2016); Bhattacharya et al. (2017); Heyl and Budich (2017). By generating the Loschmidt amplitude for initial mixed state $\rho_{0}$ via the interferometric interpretation in topological systems after a quantum quench Bhattacharya et al. (2017); Heyl and Budich (2017), it is found that the DQPT is not eliminated but rather persists at relatively high initial temperatures. This finding highlights that DQPTs possess fundamentally different characteristics compared to quantum phase transitions, which are known to disappear at finite temperatures. It is important to note that the extension of the Loschmidt amplitude to nonzero temperatures is not unique. Lang and colleagues introduced a generalized Loschmidt amplitude for mixed states Lang et al. (2018) under the parallel transport condition Sjöqvist et al. (2000); Tong et al. (2004), and similarly confirmed that DQPT is not destroyed at finite temperatures, even under sufficiently weak dissipation. The only difference between both generations of the Loschmidt amplitude is the phase factor when the state is pure and the dynamics are unitary. Notably, the Fisher zero points and rate function remain unchanged Lang et al. (2018); Tang et al. (2024). However, in our subsequent study of mixed-state DQPTs in the multi-band Kitaev model, we found that the presence of energy gaps between the system’s energy levels leads to the vanishment of DQPT at nonzero temperatures Cao et al. (2024b). This observation motivated us to further explore how different band structures influence mixed-state DQPTs.
[…]
@@
where $t_{1},\cdots,t_{m-1}$ represent intracell couplings, and $t_{m}$ denotes the intercell coupling. In contrast to the SSH model, the generalized SSH-m model reveals richer variety of topological phase transitions Anastasiadis et al. (2022); Lee et al. (2022). In the Hamiltonian (1), we have imposed the periodic boundary condition. Consequently, the Bloch Hamiltonian is obtained by
@@
In contrast, when considering the dynamics at finite temperature, the initial state is prepared as the thermal state described by a full-ranked density matrix $\rho_{0}$. The density matrix at time $t$ after the quench is given by $\rho(t)=e^{-iH(t)}\rho_{0}e^{iHt}$. Hence, a generalization of the Loschmidt amplitude to general density matrices we consider is defined as Heyl and Budich (2017); Bhattacharya et al. (2017)
@@
It is important to emphasize that the generalized Loschmidt amplitude proposed here is not limited by the parallel transport condition Sjöqvist et al. (2000); Tong et al. (2004); Lang et al. (2018); Tang et al. (2024). In our model, the initial density matrices for every $k$ take the form
@@ III DQPTs for pure states in SSH-3 model
In this section, we consider a generalization of the SSH model hosting three lattice sites in a unit cell. Notably, for the 1D SSH-3 model, the spectrum of the bulk Hamiltonian is not symmetric concerning zero Anastasiadis et al. (2022), implying that the system does not possess chiral symmetry. The winding number is also not well-defined for the SSH-3 model He and Chien (2020). However, the bulk Hamiltonian exhibits an additional interesting symmetry called point chirality, i.e. $\Gamma_{k}H_{k}\Gamma_{k}^{{\dagger}}=-H_{\pi-k}$ with $\Gamma_{k}=\mathrm{diag}(1,-1,1)$. In Fig. 1, we show the energy spectra of the bulk Hamiltonians for varying $t_{3}$ with fixed inner-cell hopping parameters $t_{1}$ and $t_{2}$. The Hamiltonian exhibits the gap closing only when $t_{1}=t_{2}=t_{3}$, which exactly corresponds to the case of the SSH model.
@@ V DQPTs in SSH-4 model
By calculating the energy spectra under different parameters, we have identified three typical energy spectrum configurations in the SSH-4 model, one of which resembles the energy spectra of the period-two Kitaev chain, as discussed in the Ref. Cao et al. (2024b). In this article, we will discuss two other cases different from the period-two Kitaev chains, with the shapes of their energy spectra shown in Fig. 8.
@@
From Fig. 8 (a), it is can be seen the system with parameters $t_{1}=t_{2}=t_{3}$ has a gap-closing point at $t_{4c}=1$. In contrast, as shown in Fig. 8 (b), the system with parameters $t_{1}=t_{3}=1$ and $t_{2}=2$ has two gap-closing point at $t_{4c}=0.5$ and $t_{4c}=2$. These critical points that collapse the energy bands will result in the topological transitions in the system, which are characterized by the change in the number of edge modes Lee et al. (2022). It is evident that based on previous experience, we can speculate that in the SSH-4 model, the behavior of the DQPT after a quench will depend on whether the quench passes through these critical points.